OpenStudy (anonymous):

1. Determine the zeros of f(x) = x3 – 3x2 – 16x + 48

OpenStudy (anonymous):

soryy it took so long @ganeshie8

ganeshie8 (ganeshie8):

aha its okay :)

ganeshie8 (ganeshie8):

\( f(x) = x^3 – 3x^2 – 16x + 48 \)

ganeshie8 (ganeshie8):

you have four terms.

ganeshie8 (ganeshie8):

look at first two terms. we can pull out x^2 ?

OpenStudy (anonymous):

yes so x²(x-3) ?

ganeshie8 (ganeshie8):

thats right ! next, look at last two terms, can we pull out -16 ?

ganeshie8 (ganeshie8):

\( f(x) = x^3 – 3x^2 – 16x + 48 \) \( = x^2(x-3) - 16(x - 3)\)

OpenStudy (anonymous):

okk , I get that part.

ganeshie8 (ganeshie8):

good :) next, see (x-3) is common

ganeshie8 (ganeshie8):

pull it out

ganeshie8 (ganeshie8):

\( f(x) = x^3 – 3x^2 – 16x + 48 \) \( = x^2(x-3) - 16(x - 3)\) \( = (x-3)(x^2 - 16)\)

OpenStudy (anonymous):

(x²-16)(x-3)

ganeshie8 (ganeshie8):

Now, to get the zeroes set it to 0

OpenStudy (anonymous):

sooo 4 and 3 ?

ganeshie8 (ganeshie8):

\((x-3)(x^2 - 16) = 0 \) \((x-3) = 0 , \ \ (x^2 - 16) = 0 \) \(x = 3 , \ \ x^2 = 16 \) \(x = 3 , \ \ x = \pm\sqrt{16} \) \(x = 3 , \ \ x = 4, \ \ \ x = -4 \)

ganeshie8 (ganeshie8):

3 zeroes, since it is a 3rd degree polynomial

OpenStudy (anonymous):

ohhh okk thankyouu veryy much, i understand it now (: . okk im going to try and do the next 3 and ill tag you in them (;

ganeshie8 (ganeshie8):

brilliant :)